can someone help me plssssssssss :)?
it takes carolyn 4 hr to type a final exam. it takes Elise 3 hr to do the same job. How long would it take them, working together, to do the typing?

Found 2 solutions by nerdybill, solver91311:
Answer by nerdybill(7384)

it takes carolyn 4 hr to type a final exam. it takes Elise 3 hr to do the same job. How long would it take them, working together, to do the typing?
.
Let x = time (hours) it takes for both
then
x(1/4 + 1/3) = 1
multiplying both sides by 12:
x(3 + 4) = 12
x(7) = 12
x = 12/7
x = 1.7 hours
or
x = 1 hour and 43 minutes

Answer by solver91311(24713)



If A can do a job in x time periods, then A can do of the job in 1 time period. Likewise, if B can do the same job in y time periods, then B can do of the job in 1 time period.

So, working together, they can do



of the job in 1 time period.

Therefore, they can do the whole job in:



time periods.


John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism